# Revisiting Speech-Lip Alignment: A Phoneme-Aware Speech Encoder for Robust Talking Head Synthesis

### v2

@@ Abstract
Speech-driven talking head synthesis tasks commonly use general acoustic features as guided speech features. However, we discovered that these features suffer from phoneme-viseme alignment ambiguity, which refers to the uncertainty and imprecision in matching phonemes with visemes.To overcome this limitation, we propose a phoneme-aware speech encoder (PASE) that explicitly enforces accurate phoneme–viseme correspondence. PASE first captures fine-grained speech and visual features, then introduces a prediction–reconstruction task to improve robustness under noise and modality absence. Furthermore, a phoneme-level alignment module guided by phoneme embeddings and contrastive learning ensures discriminative audio and visual alignment. Experimental results show that PASE achieves state-of-the-art performance in both NeRF and 3DGS rendering models. Its lip sync accuracy improves by 13.7% and 14.2% compared to the acoustic feature, producing results close to the ground truth videos.
keywords:
Multimodal Alignment , Talking Head , Speech Encoder
journal: Nuclear Physics B
[…]
1 Introduction
@@
Existing work uses general acoustic features as guided speech features. However, we discovered that these features suffer from phoneme-viseme alignment ambiguity. Ambiguity refers to the phenomenon where different phonemes (e.g., /t/ and /d/, /k/ and /g/) correspond to similar visemes (lip). As shown in Figure 1, we first decompose the phoneme–viseme alignment task into two generalizable sub-tasks: 1) extracting visual motion information from acoustic features; 2) mapping the extracted motion information to visual image sequences. As shown by the gray arrows in Figure 1, current work employs acoustic features, such as DeepSpeech [4], HuBERT [5], Wav2Vec 2.0 [6] and Whisper [7], as conditional input to the rendering model for video synthesis. However, these features are designed for tasks like speech recognition or speaker identification and focus on the discriminative power of phonemes in the acoustic domain. Consequently, phoneme-viseme alignment relies on weak alignment within the rendering model, which is simply achieved through methods like feature addition or attention mechanisms. This weak alignment does not adequately address the ambiguity between phonemes and visemes. As a result, several issues can arise when using acoustic features to drive video synthesis: 1) Inaccurate lip shape. Due to the phoneme-viseme alignment ambiguity, the alignment between the lip movements and the phonemes is inaccurate. 2) Blurry lip shape. Acoustic features do not learn the dynamic relationship between the speech signal and lip movements, resulting in a blurry lip shape.
@@
This paper proposes a universal speech encoder, PASE, which can be integrated into various talking head synthesis models to solve the issue of phoneme-viseme alignment ambiguity. PASE introduces a phoneme-aware multimodal alignment framework with contrastive learning. Specifically, PASE models the audio using an STFT spectrogram and a GRU-based encoder, which preserves high-frequency phonetic cues and captures temporal dynamics at the phoneme level. On the visual side, inspired by Wav2Lip [8] and SyncNet [9], PASE employs a multi-channel CNN to extract spatio-temporal lip features. To enhance robustness, a prediction–reconstruction task is applied, where portions of audio or visual features are randomly masked and reconstructed during training. Finally, PASE introduces a phoneme-level alignment module that explicitly enforces the audio–visual correspondence. For each phoneme, audio features serve as anchors and are fused with visual features via a cross-attention mechanism. To further enhance phoneme discrimination, phoneme embeddings are integrated into the audio features before fusion. Contrastive learning is then applied, minimizing the distance between positive audio–visual pairs while maximizing the distance to negative pairs. This design ensures that PASE focuses on phoneme-specific articulatory patterns. As shown in the orange arrows of Figure 1, PASE is designed to function independently of specific rendering models, allowing it to be seamlessly integrated into various backbone architectures to enhance the quality of synthesized faces. Experimental results show that videos synthesized using PASE outperform those synthesized using acoustic features in both NeRF and 3DGS rendering models. We summarize our contributions as follows.
  1. 1.
@@
2 Related Work
2.1 Acoustic Feature
@@
2.2 Talking Head Synthesis
In recent years, due to its applications in digital humans, virtual avatars, and video conferencing, talking head synthesis [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25], especially the real-time talking head synthesis [26, 15, 27, 17, 18, 28, 29, 30, 31, 32, 33, 19, 20, 34], has attracted significant attention. The main rendering models currently used are based on Neural Radiance Fields (NeRF) [35] and 3D Gaussian Splatting (3DGS) [36]. NeRF achieves high-fidelity scene rendering by constructing an implicit continuous volume scene representation and modeling the color and density distribution of light propagation using multi-layer perceptions. 3D Gaussian Splatting, on the other hand, parameterizes the scene geometry and appearance using dynamic Gaussian point clouds and realizes real-time dynamic rendering through differentiable rasterization. In some representative works, AD-NeRF [15], ER-NeRF [16] and TalkingGaussian [20] use DeepSpeech as the speech feature; GeneFace [18] uses HuBERT; GaussianSpeech [19] uses Wav2Vec 2.0; and Salehi et al. [37] use Whisper.
Table 1: The list of some phonemes (using consonants as an example) related to phoneme-viseme alignment ambiguity. “&” indicates that the lip shapes of the two phonemes are similar.
@@ 2.2 Talking Head Synthesis
3 Motivation
Figure 2: The lip shape diagram of some phonemes related to phoneme-viseme alignment ambiguity.
@@ 3 Motivation
3.1 Phoneme-Viseme Alignment Ambiguity
Phoneme-viseme alignment ambiguity refers to the phenomenon that different phonemes can correspond to a similar lip shape. In Table 1, we list some phonemes related to phoneme-viseme alignment ambiguity along with their corresponding pronunciation manner. Additionally, we present the specific lip shapes associated with the phoneme-viseme alignment ambiguity phenomenon in Figure 3. To explain the origin of this phenomenon in detail, we also present a vocal tract anatomical diagram in Figure 3 from Ramoo [38] to illustrate the organs that affect pronunciation. The larynx or vocal cords are the basis of pronunciation, while the lips and tongue are the articulatory organs. Although the lips are only a part of the speech production system, they are the most intuitive visual feature, which leads to the phoneme-viseme alignment ambiguity issue. For example, as shown in Figure 4, the lip shapes corresponding to phonemes /d/ and /t/ are similar. However, /d/ is a voiced consonant with strong vocal cord vibrations during pronunciation, while /t/ is a consonant with weak vocal cord vibrations. Despite the significant acoustic differences between /d/ and /t/, the speech features of /d/ and /t/ should be aligned to similar visual features in the cross-modal alignment.
3.2 Drawback of Using Acoustic Features
@@
3.3 Contrastive Learning of Speech-Lip
@@ 4.1 Overview
4.2 Audio Encoder
Most existing methods follow the traditional paradigm by directly using the Mel spectrogram, without thoroughly investigating their suitability for phoneme-viseme alignment tasks. We argue that the STFT spectrogram is a more appropriate choice. The Mel spectrogram is originally designed to mimic human auditory perception by preserving formant structures (typically within 0.3–4kHz) that convey semantic information, while suppressing high-frequency components to improve efficiency and robustness in speech recognition. However, high-frequency components (>4kHz) carry important articulatory motion cues. For instance, fricatives like /S/ and /s/ exhibit distinct energy distributions in the high-frequency. However, the high-frequency suppression in the Mel spectrogram undermines the model’s ability to capture dynamic visual cues, thereby increasing the alignment ambiguity. Furthermore, the superiority of STFT is also validated through the ablation studies 5.5.
@@ 4.2 Audio Encoder
4.3 Visual Encoder
@@
Table 2: The network architecture of the lip feature extractor.
We aim to provide accurate lip features for aligning speech and lip shapes. Inspired by works [8, 9], we designed a multi-channel CNN-based model to extract lip features. The specific architecture of this model can be found in Table 2. The model captures fine-grained spatial features of the lip images layer by layer through multiple convolutional layers. After processing the lip shape image through the CNN, an embedding is obtained with the same dimension as the speech features, representing the dynamic visual features of the lip shapes.
4.4 Prediction and Reconstruction
@@
4.5 Phoneme Level Alignment
To improve the alignment capability of the model, we propose a phoneme-level alignment strategy. Specifically, for a given phoneme $p$, we use the corresponding audio feature $A_{p}$ as the anchor and apply a cross-attention mechanism to fuse positive and negative visual features. During the fusion process, the phoneme label is embedded to obtain the phoneme label feature $F_{p}$. As formulated in Eq. [q_vector], the phoneme label feature is integrated with the audio feature to generate the query vector $Q_{p}$.
@@ 4.5 Phoneme Level Alignment
For the positive pair $(A_{p},F^{+}{fusion})$ and negative pairs $(A_{p},F^{-}{fusion})$, we employ a contrastive loss, as defined in Eq. 6, to constrain the alignment.
$\mathcal{L}_{\text{con}}=-\log\frac{\exp\left(s(A_{p},F^{+}_{fusion})/\tau\right)}{\exp\left(s(A_{p},F^{+}_{fusion})/\tau\right)+\sum\limits_{F^{-}_{fusion}}\exp\left(s(A_{p},F^{-}_{fusion})/\tau\right)}$ (6)
@@
5 Experiments
@@
Table 3: The quantitative results of video synthesis using different speech features. We highlight the best and second best results.
5.1 Experimental Settings
@@ 5.2 Quantitative Evaluation
Table 4: The quantitative results of lip accuracy. We highlight the best and second best results.
Evaluation Results. We present the quantitative results in Table 3. Our method exhibits significant advantages in lip sync accuracy. Our method's LMD surpasses the four acoustic features. In the NeRF model, our method improves LSE-C and LSE-D by 13.7% and 14.2%, respectively, compared to the HuBERT [5]. This advantage is further amplified in the 3DGS model. Notably, the LSE-C and LSE-D of our method are close to the ground truth video, indicating the fine-grained capture of speech dynamic information by the STFT-GRU combination and the effectiveness of the cross-modal alignment framework. Furthermore, our approach outperforms state-of-the-art face synthesis methods, including AVHubert [41], Wav2Lip [8], SyncTalk [17], and TalkingGaussian [20]. It is worth noting that SyncTalk [17] employs specially designed speech features, while TalkingGaussian [20] relies on DeepSpeech [4]. We also used out-of-distribution speech to drive video synthesis in the NeRF model. We present the experimental results in Table 4. We used LMD, LSE-C, and LSE-D as metrics. The experimental results show that our method outperforms the four acoustic features in terms of lip sync accuracy.
@@
Table 5: The results of the user study. The rating is on a scale of 1-5. A higher scores indicate better performance. We highlight the best and second best results.
5.3 Qualitative Evaluation
@@
User Study. We conducted a user study to assess the quality of the synthesized videos effectively. We sampled 20 video clips from the quantitative evaluation and invited 12 volunteers to participate in the study. We used the mean opinion score (MOS) as the metric. Volunteers were asked to rate the synthesized videos on three aspects: 1) Lip-sync Accuracy, 2) Image Quality, and 3) Video Realness. The average scores for each method are presented in Table 5. As shown, our method significantly outperforms the comparison methods in terms of Lip-sync Accuracy, indicating the effectiveness of the cross-modal alignment framework.
5.4 Model Inference Efficiency
We present the inference efficiency of different models in Table 6. The experimental results show that the proposed PASE achieves superior hardware efficiency while maintaining competitive inference speed. Specifically, PASE attains a 8.8× parameter reduction compared to Wav2vec 2.0 (37.7M vs 317M) with only 38.4% longer inference time (4.72s vs 3.41s), while significantly reducing GPU memory consumption by 57.3% (2051.5MB vs 4805.3MB). Compared to the lightweight Whisper, PASE shows comparable computational efficiency (4.72s vs 5.03s). These metrics collectively verify our design's effectiveness in balancing computational efficiency and phoneme-viseme alignment accuracy.
@@ 5.4 Model Inference Efficiency
Table 6: The results of the inference efficiency.
5.5 Ablation Study
@@
Table 7: The results of ablation study on speech spectrogram and modeling. We highlight the best and second best results.
To evaluate the effectiveness of the PASE framework, we conducted ablation experiments from two dimensions: speech spectrograms and modeling. The experimental results are presented in Table 7. As shown, the combination of STFT spectrogram and GRU significantly outperforms other variants in both video fidelity and lip sync accuracy, highlighting the necessity of the collaborative design of modules in PASE.
@@
6 Conclusion
To address the challenge of phoneme–viseme alignment ambiguity, we introduce PASE, a universal speech encoder that can be flexibly integrated into diverse talking portrait synthesis models. Unlike conventional approaches that align linguistic representations, PASE enforces phoneme-level cross-modal alignment between speech and lip shapes, guided by phoneme embeddings and contrastive learning. Furthermore, by combining STFT spectrograms with a GRU-based encoder, PASE effectively captures fine-grained temporal–frequency features. This design enhances robustness against noise and modality absence, while providing more discriminative and generalizable representations to support subsequent rendering models. Experimental results show that PASE achieves state-of-the-art performance on both NeRF and 3DGS rendering models. Notably, in terms of lip sync accuracy, PASE improves LSE-C and LSE-D by 13.7% and 14.2%, compared to the acoustic feature, and produces results close to the ground truth videos. Ablation experiments further present the effectiveness of the STFT-GRU combination.
7 Acknowledgements
This work is supported by the Natural Science Foundation of China (NSFC) under the grants No. 62572358, 621723064, and 62372334.

### v3

# PASE: Phoneme-Aware Speech Encoder to Improve Lip Sync Accuracy for Talking Head Synthesis

@@ Abstract
Recent talking head synthesis works typically adopt speech features extracted from large-scale pre-trained acoustic models. However, the intrinsic many-to-many relationship between speech and lip motion causes phoneme–viseme alignment ambiguity, leading to inaccurate and unstable lips. To further improve lip sync accuracy, we propose PASE (Phoneme-Aware Speech Encoder), a novel speech representation model that bridges the gap between phonemes and visemes. PASE explicitly introduces phoneme embeddings as alignment anchors and employs a contrastive alignment module to enhance the discriminability between corresponding audio–visual pairs. In addition, a prediction and reconstruction task is designed to improve robustness under noise and partial modality absence. Experimental results show PASE significantly improves lip sync accuracy and achieves state-of-the-art performance across both NeRF- and 3DGS-based rendering frameworks, outperforming conventional methods based on acoustic features by 13.7% and 14.2%, respectively. Importantly, PASE can be seamlessly integrated into diverse talking head pipelines to improve the lip sync accuracy without architectural modifications.
I Introduction
@@
Existing work uses general acoustic features as guided speech features. However, we discovered that these features suffer from phoneme-viseme alignment ambiguity. Ambiguity refers to the phenomenon where different phonemes (e.g., /t/ and /d/, /k/ and /g/) correspond to similar visemes (lip). As shown in Figure 1, we first decompose the phoneme–viseme alignment task into two generalizable sub-tasks: 1) extracting visual motion information from acoustic features; 2) mapping the extracted motion information to visual image sequences. As shown by the upper part in Figure 1, current work employs acoustic features, such as DeepSpeech [4], HuBERT [5], Wav2Vec 2.0 [6] and Whisper [7], as conditional input to the rendering model for video synthesis. However, these features are designed for tasks like speech recognition or speaker identification and focus on the discriminative power of phonemes in the acoustic domain. Consequently, phoneme-viseme alignment relies on weak alignment within the rendering model, which is simply achieved through methods like feature addition or attention mechanisms. This weak alignment does not adequately address the ambiguity between phonemes and visemes. As a result, several issues can arise when using acoustic features to drive video synthesis: 1) Inaccurate lip shape. Due to the phoneme-viseme alignment ambiguity, the alignment between the lip movements and the phonemes is inaccurate. 2) Blurry lip shape. Acoustic features do not learn the dynamic relationship between the speech signal and lip movements, resulting in a blurry lip shape.
@@
This paper proposes a universal speech encoder, PASE, which can be integrated into various talking head synthesis models to solve the issue of phoneme-viseme alignment ambiguity. PASE introduces a phoneme-aware multimodal alignment framework with contrastive learning. Specifically, PASE models the audio using an STFT spectrogram and a GRU-based encoder, which preserves high-frequency phonetic cues and captures temporal dynamics at the phoneme level. On the visual side, inspired by Wav2Lip [8] and SyncNet [9], PASE employs a multi-channel CNN to extract spatio-temporal lip features. To enhance robustness, a prediction–reconstruction task is applied, where portions of audio or visual features are randomly masked and reconstructed during training. Finally, PASE introduces a phoneme-level alignment module that explicitly enforces the audio–visual correspondence. For each phoneme, audio features serve as anchors and are fused with visual features via a cross-attention mechanism. To further enhance phoneme discrimination, phoneme embeddings are integrated into the audio features before fusion. Contrastive learning is then applied, minimizing the distance between positive audio–visual pairs while maximizing the distance to negative pairs. This design ensures that PASE focuses on phoneme-specific articulatory patterns. As illustrated in the lower part of Fig. 1, PASE leverages the semantic layer (phonemes) shared between the speech and visual modalities as alignment anchors. This design exploits not only temporal relationship constraints but also the stable semantic structure, thereby enhancing the consistency of cross-modal alignment. Experimental results show that videos synthesized using PASE outperform those synthesized using acoustic features in both NeRF and 3DGS rendering models. We summarize our contributions as follows.
  1. 1)
@@
II Related Work
II-A Acoustic Feature
@@
II-B Talking Head Synthesis
In recent years, due to its applications in digital humans, virtual avatars, and video conferencing, talking head synthesis [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25], especially the real-time talking head synthesis [26, 15, 27, 17, 18, 28, 29, 30, 31, 32, 33, 19, 20], has attracted significant attention. The main rendering models currently used are based on Neural Radiance Fields (NeRF) [34] and 3D Gaussian Splatting (3DGS) [35]. NeRF achieves high-fidelity scene rendering by constructing an implicit continuous volume scene representation and modeling the color and density distribution of light propagation using multi-layer perceptions. 3D Gaussian Splatting, on the other hand, parameterizes the scene geometry and appearance using dynamic Gaussian point clouds and realizes real-time dynamic rendering through differentiable rasterization. In some representative works, AD-NeRF [15], ER-NeRF [16] and TalkingGaussian [20] use DeepSpeech as the speech feature; GeneFace [18] uses HuBERT; GaussianSpeech [19] uses Wav2Vec 2.0; and Salehi et al. [36] use Whisper.
TABLE I: The list of some phonemes (using consonants as an example) related to phoneme-viseme alignment ambiguity. “&” indicates that the lip shapes of the two phonemes are similar.
@@ II-B Talking Head Synthesis
III Motivation
Figure 2: The illustration of the many-to-many relationship between speech and visual modalities. Similar lip shapes correspond to different phonemes, while the same phoneme corresponds to different lip shapes under different emotions.
@@ III Motivation
III-A Phoneme-Viseme Alignment Ambiguity
As shown in Figure 2, the relationship between speech and visual modalities is inherently many-to-many. For certain types of phonemes (e.g., plosives), similar lip shapes may correspond to different phonemes. Moreover, the same phoneme can produce inconsistent lip shapes under different emotional states. This many-to-many relationship results in phoneme–viseme alignment ambiguity. In Table I, we list some phonemes related to phoneme-viseme alignment ambiguity along with their corresponding pronunciation manner. To explain the origin of this phenomenon in detail, we also present a vocal tract anatomical diagram in Figure 3 from Ramoo [37] to illustrate the organs that affect pronunciation. The larynx or vocal cords are the basis of pronunciation, while the lips and tongue are the articulatory organs. Although the lips are only a part of the speech production system, they are the most intuitive visual feature, which leads to the phoneme-viseme alignment ambiguity issue. For example, as shown in Figure 4, the lip shapes corresponding to phonemes /d/ and /t/ are similar. However, /d/ is a voiced consonant with strong vocal cord vibrations during pronunciation, while /t/ is a consonant with weak vocal cord vibrations. Despite the significant acoustic differences between /d/ and /t/, the speech features of /d/ and /t/ should be aligned to similar visual features in the cross-modal alignment.
III-B Drawback of Using Acoustic Features
@@
III-C Contrastive Learning of Speech-Lip
@@ IV-A Overview
IV-B Audio Encoder
Most existing methods follow the traditional paradigm by directly using the Mel spectrogram, without thoroughly investigating their suitability for phoneme-viseme alignment tasks. We argue that the STFT spectrogram is a more appropriate choice. The Mel spectrogram is originally designed to mimic human auditory perception by preserving formant structures (typically within 0.3–4kHz) that convey semantic information, while suppressing high-frequency components to improve efficiency and robustness in speech recognition. However, high-frequency components (>4kHz) carry important articulatory motion cues. For instance, fricatives like /S/ and /s/ exhibit distinct energy distributions in the high-frequency. However, the high-frequency suppression in the Mel spectrogram undermines the model’s ability to capture dynamic visual cues, thereby increasing the alignment ambiguity. Furthermore, the superiority of STFT is also validated through the ablation studies V-E.
@@ IV-B Audio Encoder
IV-C Visual Encoder
TABLE II: The network architecture of the lip feature extractor.
@@ IV-C Visual Encoder
We aim to provide accurate lip features for aligning speech and lip shapes. Inspired by works [8, 9], we designed a multi-channel CNN-based model to extract lip features. The specific architecture of this model can be found in Table II. The model captures fine-grained spatial features of the lip images layer by layer through multiple convolutional layers. After processing the lip shape image through the CNN, an embedding is obtained with the same dimension as the speech features, representing the dynamic visual features of the lip shapes.
IV-D Prediction and Reconstruction
@@
IV-E Phoneme Level Alignment
To improve the alignment capability of the model, we propose a phoneme-level alignment strategy. Specifically, for a given phoneme $p$, we use the corresponding audio feature $A_{p}$ as the anchor and apply a cross-attention mechanism to fuse positive and negative visual features. During the fusion process, the phoneme label is embedded to obtain the phoneme label feature $F_{p}$. As formulated in Eq. 4, the phoneme label feature is integrated with the audio feature to generate the query vector $Q_{p}$.
@@ IV-E Phoneme Level Alignment
[…]
$\mathcal{L}_{\text{con}}=-\log\frac{\exp\left(s(A_{p},F^{+}_{fusion})/\tau\right)}{\exp\left(s(A_{p},F^{+}_{fusion})/\tau\right)+\sum\limits_{F^{-}_{fusion}}\exp\left(s(A_{p},F^{-}_{fusion})/\tau\right)}$
@@
V Experiments
TABLE III: The quantitative results of video synthesis using different speech features. We highlight the best and second best results.
@@ V Experiments
V-A Experimental Settings
@@ V-B Quantitative Evaluation
TABLE IV: The quantitative results of lip accuracy. We highlight the best and second best results.
@@
Evaluation Results. We present the quantitative results in Table III. Our method exhibits significant advantages in lip sync accuracy. Our method's LMD surpasses the four acoustic features. In the NeRF model, our method improves LSE-C and LSE-D by 13.7% and 14.2%, respectively, compared to the HuBERT [5]. This advantage is further amplified in the 3DGS model. Notably, the LSE-C and LSE-D of our method are close to the ground truth video, indicating the fine-grained capture of speech dynamic information by the STFT-GRU combination and the effectiveness of the cross-modal alignment framework. Furthermore, our approach outperforms state-of-the-art face synthesis methods, including AVHubert [40], Wav2Lip [8], SyncTalk [17], and TalkingGaussian [20]. It is worth noting that SyncTalk [17] employs specially designed speech features, while TalkingGaussian [20] relies on DeepSpeech [4]. We also used out-of-distribution speech to drive video synthesis in the NeRF model. We present the experimental results in Table IV. We used LMD, LSE-C, and LSE-D as metrics. The experimental results show that our method outperforms the four acoustic features in terms of lip sync accuracy.
TABLE V: The results of the user study. The rating is on a scale of 1-5. A higher scores indicate better performance. We highlight the best and second best results.
@@
V-C Qualitative Evaluation
@@
User Study. We conducted a user study to assess the quality of the synthesized videos effectively. We sampled 20 video clips from the quantitative evaluation and invited 12 volunteers to participate in the study. We used the mean opinion score (MOS) as the metric. Volunteers were asked to rate the synthesized videos on three aspects: 1) Lip-sync Accuracy, 2) Image Quality, and 3) Video Realness. The average scores for each method are presented in Table V. As shown, our method significantly outperforms the comparison methods in terms of Lip-sync Accuracy, indicating the effectiveness of the cross-modal alignment framework.
V-D Model Inference Efficiency
We present the inference efficiency of different models in Table VI. The experimental results show that the proposed PASE achieves superior hardware efficiency while maintaining competitive inference speed. Specifically, PASE attains a 8.8× parameter reduction compared to Wav2vec 2.0 (37.7M vs 317M) with only 38.4% longer inference time (4.72s vs 3.41s), while significantly reducing GPU memory consumption by 57.3% (2051.5MB vs 4805.3MB). Compared to the lightweight Whisper, PASE shows comparable computational efficiency (4.72s vs 5.03s). These metrics collectively verify our design's effectiveness in balancing computational efficiency and phoneme-viseme alignment accuracy.
TABLE VI: The results of the inference efficiency.
@@ V-D Model Inference Efficiency
V-E Ablation Study
TABLE VII: The results of ablation study on speech spectrogram and modeling. We highlight the best and second best results.
@@ V-E Ablation Study
To evaluate the effectiveness of the PASE framework, we conducted ablation experiments from two dimensions: speech spectrograms and modeling. The experimental results are presented in Table VII. As shown, the combination of STFT spectrogram and GRU significantly outperforms other variants in both video fidelity and lip sync accuracy, highlighting the necessity of the collaborative design of modules in PASE.
@@
VI Conclusion
In this paper, we addressed the challenge of phoneme–viseme alignment ambiguity in talking head synthesis. To this end, we proposed PASE (Phoneme-Aware Speech Encoder), a novel and generalizable speech representation model that explicitly bridges the gap between phonemes and visemes. By introducing phoneme embeddings as alignment anchors and incorporating a contrastive alignment module, PASE effectively enhances the discriminability and consistency of audio–visual correspondence. Moreover, the prediction and reconstruction task further improves robustness under noisy environments and partial modality absence. Comprehensive experiments on both NeRF- and 3DGS-based rendering frameworks demonstrate that PASE significantly improves lip sync accuracy, achieving 13.7% and 14.2% gains over conventional acoustic features. Furthermore, PASE serves as a plug-and-play speech encoder, seamlessly integrated into existing talking head pipelines to improve the lip sync accuracy without architectural modifications.